OpenStudy (anonymous):

Sally gave her brother $3.95 in quarters and dimes. If the number of quarters is 6 more than the number of dimes, how many quarters did Sally give her brother

OpenStudy (mathstudent55):

First, choose a variable to represent the number of dimes and another variable to represent the number of quarters..

OpenStudy (mathstudent55):

How about let d = number of dimes and let q = number of quarters

OpenStudy (mathstudent55):

Ok so far?

OpenStudy (mathstudent55):

@daniellelynn Are you following?

OpenStudy (anonymous):

no to be honest. this is my biggest struggle and im having the worst time understanding

OpenStudy (mathstudent55):

To solve this problem you need a system of equations. A system of equation is 2 or more equations solved together to arrive at an answer. The first step we need is to let two letters, called variables, represent what we are looking for.

OpenStudy (mathstudent55):

We could use x and y for the variables, but to make it easy to remember what the variables represent, we can let the variable d equal the number of dimes and the variable q equal the number of quarters.

OpenStudy (mathstudent55):

From this statement: "the number of quarters is 6 more than the number of dimes" we get: q = d + 6 This is our first equation.

OpenStudy (mathstudent55):

Now we use the value of the coins and the total to find another equation.

OpenStudy (mathstudent55):

A dime is worth $0.10 A quarter is worth $0.25

OpenStudy (mathstudent55):

d dimes are worth 0.1d q quarters are worth 0.25q

OpenStudy (mathstudent55):

The problem states: "$3.95 in quarters and dimes" That means 0.1d + 0.25q = 3.95 That is our second equation.

OpenStudy (mathstudent55):

Here are the two equations: q = d + 6 0.1d + 0.25q = 3.95

OpenStudy (mathmate):

@daniellelynn oohoo! @mathstudent55 has been writing for 15 minutes and he would appreciate some response... please!

OpenStudy (mathstudent55):

Since the first equation is already solved for q, we can use substitution. We replace q of the second equation with what q is equal to in the first equation, d + 6. 0.1d + 0.25(d + 6) = 3.95 Distribute the 0.25: 0.1d + 0.25d + 1.5 = 3.95 Add 0.1d and 0.25d together. Subtract 1.5 from both sides: 0.35d = 2.45 d = 7 There are 7 dimes. Since the number of quarters is 6 more than dimes, there are 7 + 6 = 13, 13 quarters.

OpenStudy (mathmate):

Another method to solve this problem: You can always set up a system of linear equations with 2 unknowns and then solve the system, as @mathstudent55 has clearly explained. However, if you want to do it in your head, this is what you could do. 1. There are \(6~~more~~quarters\) than dimes, so if we take away the 6 quarters, we can subtract 6*$0.25=$1.50 from the total, we are left with $2.45 with equal number of quarters and dimes. 2. Divide $2.45 by (0.25+0.10)=0.35 gives 7. 3. So there are 7 dimes and (7+6)=13 quarters.

OpenStudy (anonymous):

thank you for the help! thank you so much!!